OpenStudy (anonymous):

Hi, I'm trying to solve a taylor series problem (actually, a maclaurin series in this case) for.. f(x)=ln(1-2x) I need to find the interval of convergence and the radius of convergence. I don't know what to do past the step where I write down the series itself. ...ln(1-2x)= 0 + -2(x-0)/1! + -4(x-0)^2/2! + -16(x-0)^3/3! + ...

OpenStudy (anonymous):

how did you get that series?

OpenStudy (anonymous):

For the function f(x)=ln(1-2x), I found out the first few derivatives of f(x), then the values when x=0. Then, I wrote these values in the ''pattern'' for the maclaurin series. And that's where I'm stuck...

OpenStudy (anonymous):

These where my first derivatives: f(x)=ln(1-2x) f'(x)=-2/(1-2x) f''(x)=-4/(1-2x)^2 f'''(x)=-16/(1-2x)^2 and it keeps going... f(0)=0 f'(0)=-2 f''(0)=-4 f'''(0)=-16... These are the values

OpenStudy (anonymous):

looks good to me. it would be prettier if you wrote \[x\] instead of \[x-0\] for each term

OpenStudy (anonymous):

Yes, I know. I just keep the 0's to remind me of where ''a'' goes in a taylor series form.

OpenStudy (anonymous):

1/x-1 = Sum of x^n from 0 to infinity 1/(2x-1) = \[\sum _{n=0}^{\infty } (2x)^n\] integrate that

OpenStudy (anonymous):

1+ 2x + 4x^2 +8 x^3 integrate into x + 2/2x^2+ 4/3 x^3 + 8/4 x^4 =

OpenStudy (anonymous):

which is \[\sum _{n=1}^{\infty } \frac{2^{n-1}}{n} x^n\]=\[-\frac{1}{2} \text{Log}[1-2 x]\]

OpenStudy (anonymous):

\[-2\sum _{n=1}^{\infty } \frac{2^{n-1}}{n} x^n\]= ln(1-2x)

OpenStudy (anonymous):

Ok, thanks for the help ! I'll write these down and try to understand them.

OpenStudy (anonymous):

Integral of 1/(2x-1) is −1/2Log[1−2x]